OpenStudy (anonymous):

Let f(x) be the function \frac{1}{x+8}. Then the quotient \frac{f(5+h)-f(5)}{h} can be simplified to \frac{-1}{ah+b} for: a= and b=

OpenStudy (anonymous):

\[\frac{1}{x+8}\]

OpenStudy (anonymous):

\[\frac{f(5+h)-f(5)}{h} =\frac{-1}{ah+b}\]

OpenStudy (anonymous):

you should get \[\frac{-1}{(5+8)(5+h+8)}\]

OpenStudy (anonymous):

\[\frac{-1}{13(13+h)}\] \[\frac{-1}{169+13h}\] i think

OpenStudy (anonymous):

thats correct

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw